### v1

# KPNDepth: Depth Estimation of Lane Images under Complex Rainy Environment

Abstract
With the development of deep neural network generative models in recent years, significant progress has been made in the research of depth estimation in lane scenes. However, current research achievements are mainly focused on clear daytime scenarios. In complex rainy environments, the influence of rain streaks and local fog effects often leads to erroneous increases in the overall depth estimation values in images. Moreover, these natural factors can introduce disturbances to the accurate prediction of depth boundaries in images. In this paper, we investigate lane depth estimation in complex rainy environments. Based on the concept of convolutional kernel prediction, we propose a dual-layer pixel-wise convolutional kernel prediction network trained on offline data. By predicting two sets of independent convolutional kernels for the target image, we restore the depth information loss caused by complex environmental factors and address the issue of rain streak artifacts generated by a single convolutional kernel set. Furthermore, considering the lack of real rainy lane data currently available, we introduce an image synthesis algorithm, RCFLane, which comprehensively considers the darkening of the environment due to rainfall and local fog effects. We create a synthetic dataset containing 820 experimental images, which we refer to as RainKITTI, on the commonly used depth estimation dataset KITTI. Extensive experiments demonstrate that our proposed depth estimation framework achieves favorable results in highly complex lane rainy environments.
Index Terms:
depth estimation, complex rainy environment, convolution kernel prediction, image reconstruction
I Introduction
Depth estimation is one of the key tasks in the field of computer vision, aiming to infer distance information from each pixel in an image to the camera. Based on this depth information, the geometric layout of a three-dimensional scene can be quickly understood, providing crucial visual information support for areas such as autonomous driving and intelligent transportation systems. Traditional depth estimation often relies on SLAM methods [1], which, in the absence of prior knowledge, use sensors (such as high-precision depth cameras, LiDAR, etc.) for self-localization and environmental modeling. However, high-precision depth cameras and LiDAR systems often come with high equipment costs. To address this issue, in recent years, with the advancement of deep neural networks, estimating pixel-level dense depth from a single image captured by a regular camera is becoming feasible [2]. As research progresses, various types of deep learning-based monocular depth estimation frameworks have been proposed, including supervised learning methods that utilize depth information from 3D sensor scans as training ground truth [3], as well as self-supervised and unsupervised methods based on binocular epipolar geometric constraints [4] and monocular frame sequence pose estimation [5], separately. These methods have garnered widespread attention in recent years due to their low cost and ability to generalize across various scenarios.
However, current research on monocular depth estimation has long faced the challenge of complex weather conditions. Additional natural factors such as occlusion by rain streaks and local fog effects [6] can lead to errors and loss of depth information in images. Moreover, the KITTI dataset [7], which is widely used for monocular depth estimation, only contains road images during clear daytime and lacks driving data on rainy days and nights. Therefore, existing studies [8][9] rely on synthetic data for training. However, existing research indicates that unsupervised approaches using geometric constraints instead of ground truth face an inevitable dilemma when confronted with augmented datasets. Existing depth estimation algorithms, especially self-supervised and unsupervised methods, are based on the assumption of photometric consistency [10], which is generally satisfied in daytime images where the sun serves as the primary light source. However, in complex rainy conditions, multiple light sources such as streetlights, car lights, and high-reflectance surfaces due to rainwater violate inter-frame photometric consistency, leading to significant errors in photometric loss calculation and severely impacting the estimation of depth information in images.
Therefore, stemming from the aforementioned challenges, we aim to propose a new research approach to reduce the strong dependence of depth estimation research in complex rainy environments on photometric consistency.
When introducing complex rainy lane images, we diverge from current mainstream approaches [11][12] that directly input synthetic images alongside original images into the depth estimation framework for self-supervision. Instead, we decompose depth estimation in complex environments into two subtasks: image reconstruction and depth estimation. By constructing an efficient and accurate image reconstruction network, DLKPN, trained on offline data, we aim to restore depth information in synthetic images fed into the depth estimation network. This approach aims to reduce the reliance of the proposed Added Depth Estimation Model of DLKPN (KPNDepth) framework on the assumptions of photometric consistency and static scene conditions in input images. Our main contributions are as follows:
  • Taking into account the environmental darkening and local fog effects caused by rain, we have proposed an algorithm called RCFLane for synthesizing complex rainy lane images. Based on this algorithm, we have synthesized RainKITTI using the KITTI dataset [7], and we plan to release this dataset.
  • Building upon the concept of convolutional kernel prediction [13], we have developed the image reconstruction network DLKPN to address rain streak artifacts in complex rainy lane image reconstruction. We have achieved efficient registration of input images on RainKITTI.
  • ·We have tested our model on RainKITTI, achieving more effective image reconstruction compared to existing methods [14][15]. Furthermore, the proposed KPNDepth framework has shown more accurate depth estimation results compared to baseline models [16].
II PROPOSED METHOD
In this section, we will introduce the complete KPNDepth framework, the RCFLane algorithm for synthesizing real rainy lane images, and the DLKPN as the core of this framework.
II-A RCFLane: Building Real Lane Data in Rainy Cloudy and Foggy Days
Due to the effect of rain on ambient lighting and camera exposure, it is almost impossible to take a realistic set of photos from the same location, leaving the only difference between the two photos being the presence or absence of rain streaks. Most existing commonly used datasets add 2D rain streaks to original image data. However, for the task we are studying, these datasets only consider the factor of rain streaks when synthesizing rainy day datasets, while ignoring the darkening of the environment and local fog effects caused by rainfall. These factors are crucial in our depth estimation task and cannot be overlooked. Although the RainCityscapes dataset proposed in [17] introduces fog when synthesizing data, it focuses on urban landscapes rather than lane data, which does not align with the training purpose of our model. Therefore, this paper will start with the commonly used KITTI dataset for lane depth estimation, aiming to simulate realistic rain characteristics as much as possible on it to construct subsequent training and testing data.
In general, factors such as rain and fog can be considered a form of degradation, and it is reasonable to use image filtering methods to address them. Specifically, for the original lane images $O(x)\in R^{I\times J}$, based on existing research, we propose a more efficient rainy day image synthesis strategy called RCFLane:
$O_{1}(x)=\alpha O(x)+\beta R(x)$ (1)
$O_{2}(x)=\gamma O_{1}(x)+(1-\gamma)D(x)$ (2)
$\hat{O}(x)=O_{2}(x)\cdot td(x)+A(1-td(x))$ (3)
where $O_{1}(x)$ and $O_{2}(x)$ represent intermediate images in the synthetic data generation process, $\hat{O}(x)$ represents the final synthesized rainy image, $x$ denotes the two-dimensional pixel coordinates, $R(x)$ stands for the rain layer, $D(x)$ signifies the mask layer, $tr(x)$ represents the transmission layer, and $A$ represents the atmospheric light, $\alpha$, $\beta$ and $\gamma$ are used to control the retention weights of the original image, the introduction weight of the rain layer, and the introduction weight of the mask layer, respectively. For specific details, please refer to Fig. 1 for the detailed process.
Figure 1: RCFLane Algorithm Implementation Example.
II-A1 Rain Layer
In the rain operation of RCFLane, we rely on a Gaussian blur kernel. By introducing three parameters - rotation center coordinates, rotation angle, and scale factor - we construct an affine transformation matrix to perform image rain filtering by rotating and generating a Gaussian blur kernel based on the unit focus matrix.
It is important to note that the rain layer constructed using Gaussian blur has a black background. We only want to overlay the white rain lines, so in Eq.(1) we set the parameter $\alpha$
$\alpha=\frac{255-O(x)}{255}$ (4)
where $O(x)$ contains RGB pixel information across three channels. For the black parts of the rain layer, when $\alpha=1$ all original image information is retained. Another parameter $\beta$ is manually determined and controls the introduction weight of the rain layer.
II-A2 Mask Layer
In reality, the impact of rain on images is not only manifested in the obscuration caused by rain streaks but also in the deterioration of ambient lighting conditions. To reflect the changes in lighting conditions due to rainy days, this paper introduces a mask layer on top of the previous $O_{1}(x)$ in Eq.(2). The mask layer can be understood as a fully black curtain, weighted by a parameter $\gamma$ set by humans, to simulate environmental lighting from day to night. This allows for a more realistic simulation of lighting conditions on rainy days, making the synthetic data closer to real rainy day lighting situations.
II-A3 Transmission Layer
The impact of rain curtains on images presents a certain fog-like characteristic on the images visually due to the scene depth. This fog-like characteristic becomes more pronounced as the distance from the camera increases.
In this paper, the addition of fog in Eq.(3) is based on the standard optical model. In imaging under hazy weather conditions, the reduction of reflected energy causes light transmission attenuation, and atmospheric light scatters around, enhancing the surrounding brightness in foggy conditions. Therefore, the pixel value at a certain point in the foggy image consists of two parts
$td(x)=e^{-\lambda d(x)}$ (5)
where $\lambda$ is the attenuation coefficient to control the thickness of the fog, and here $d(x)$
$d(x)=-|x-x_{mid}|+S$ (6)
The parameter $S$ represents the fog scale, where there is a negative correlation between $d(x)$ (the distance from a pixel point to the center point) and the fog scale. The reason why the RCFLane algorithm fixes the fogging center at the image center instead of incorporating image depth information into fog synthesis, as suggested in [6], is twofold. Firstly, the method proposed in [6] demands precise image depth values, we hope to explore a less restrictive lane rainy day image synthesis method. Secondly, and more importantly, for lane data, the center of the image often represents the end of the road extension, that is, the farthest distance from the camera. Based on the nature of rain curtain fog being distance-dependent, we can infer that the fog thickness reaches its maximum at the image center and decreases towards the periphery. Therefore, fixing the fogging center at the image center is deemed reasonable.
Fig. 2 shows the intermediate image obtained by the above three steps.
Figure 2: Example of Intermediate Image Generated by RCFLane Algorithm.
II-B Dual-layer Pixel-wise Convolution Kernel Prediction Network
Whether it is rain or the decrease in illumination caused by rain curtains and local fog effects can be considered as image degradation. We hope to effectively restore these degradations before inputting the images into the depth estimation network, thereby addressing the depth increase and local pseudo-artifacts caused by rain curtains. Additionally, due to downstream requirements, today’s depth estimation tasks often have high requirements for real-time prediction, which reminds us to pay attention to the inference time required for the reconstruction algorithm.
Inspired by recent research on image restoration and rain removal [13][18][19], we aim to establish a kernel prediction network for rain removal, predicting convolutional kernels instead of directly predicting the pixel values of the target pixels. It has been proposed in [13][18][19] that predicting pixel-wise convolutional kernels for image restoration offers advantages in terms of convergence speed and inference speed, aligning well with the objectives of our study.
The KPN network operates on each pixel $p$ to be denoised in the synthesized rainy image $\hat{O}(x)\in R^{I\times J}$
$\tilde{O}(p)=W(p)\ast K_{p}$ (7)
where $\tilde{O}(p)$ represents the pixel value at point $p$ in the reconstructed image, $W(p)\in R^{K\times K}$ represents a $K\times K$ window centered at $p$ in $\hat{O}(x)$, $\ast$ denotes pixel-wise filtering operation, and represents the pixel-wise convolutional kernel predicted for pixel $p$.
Figure 3: Display of Prediction Results of Single-layer KPN on RainKITTI.
Figure 4: DLKPN Structure Diagram.
However, through experiments, we found that the degradation in $\hat{O}(x)$ compared to $O(x)$ is not solely due to rain streaks. Introducing a single KPN network with pixel-wise convolutional kernels predicted for each pixel cannot fully explain this degradation. As a result, the reconstructed image still retains some degradation information, manifested notably as rain streak artifacts. Fig. 3 specifically demonstrates this issue using predictions on RainKITTI data. Detailed experimental results will be presented in Section IV. Therefore, we have constructed a Dual-layer Pixel-wise Convolution Kernel Prediction Network (DLKPN). This network is trained offline on two independent sets of data. One layer is dedicated to recovering most of the degradation information, while the other layer focuses primarily on removing the artifacts caused by rain streaks. For each pixel $p$ in $\hat{O}(x)$, the network will predict two sets of independent pixel-wise convolution kernels $K_{p1}$ and $K_{p2}$
$\tilde{O}_{m}(p)=W(p)\ast K_{p1}$ (8)
$\tilde{O}_{f}(p)=W(\tilde{O}_{m}(p))\ast K_{p2}$ (9)
where $\tilde{O}_{m}(p)$ and $\tilde{O}_{f}(p)$ represent the predicted values of pixel $p$ in the first and second layers of the network, respectively. In practical computation, we adopt the hierarchical dilation strategy from [18]. We expand the pixel-wise convolution kernels $K_{p1}$ and $K_{p2}$ using a coefficient $r=0,1,2,3$ to construct dilated convolution kernels, which are aimed at filtering out rain streaks at multiple scales. Fig. 4 illustrates the specific structure of DLKPN.
II-C Added Depth Estimation Model of DLKPN
The DLKPN model proposed in this paper addresses the reconstruction of rainy lane images, with the advantage of fast inference speed, aligning well with the real-time requirements of depth estimation tasks. We will offline train the required pre-trained DLKPN model on the RainKITTI dataset and integrate it with traditional depth estimation models to enhance the accuracy of depth estimation in rainy environments. For ease of description, we will refer to this framework as KPNDepth.
As a baseline, we have chosen LapDepth[16], a model that utilizes Laplacian pyramids as part of the decoder structure. It takes the encoded features and feeds them into different streams, with each stream responsible for decoding depth residuals. By decomposing the Laplacian pyramid to define depth residuals and gradually combining the corresponding outputs from coarse to fine, it reconstructs the final depth image. The overall structure of the KPNDepth framework can be referenced in Fig. 5.
Figure 5: The Overall Structure of The KPNDepth Framework.
III EXPERIMENTS
III-A Datasets
Since this paper focuses on rainy lane depth estimation and there is currently a lack of publicly available rainy datasets specialized in lane information, we started from the commonly used public depth estimation dataset KITTI[7] and used the RCFLane algorithm to synthesize the RainKITTI dataset used in this study. The KITTI dataset is one of the most commonly used benchmark datasets for computer vision algorithms in autonomous driving scenarios internationally, containing real image data from urban, rural, and highway scenes. We randomly selected 820 images from the left camera images of the $2011\_09\_28$ and $2011\_09\_29$ sequences in KITTI and synthesized RainKITTI through rain image synthesis. The resolution of the synthesized images will be kept consistent with the original images (1238 x 374). For each synthesized image, we retained its ground truth depth, which will be used to evaluate the performance of our depth estimation model.
  • Training Sets:For the RainKITTI dataset, we have partitioned 715 images as the training set to train the DLKPN network structure.
  • Testing Sets:In this experiment, we have set aside 105 images from RainKITTI as the testing set. This testing set will be used to evaluate the performance of the dual-layer KPN network reconstruction. For the reconstructed images, we will assess the specific performance of the KPNDepth framework in rainy lane depth estimation based on their real depth information and depth network predictions.
III-B Experimental Setup
All our experiments and inferences were conducted on a single NVIDIA GeForce RTX 4090 GPU using the PyTorch framework. The network architecture of the depth estimation model remains consistent with LapDepth [16].
When using the RCFLane algorithm to synthesize rainy images, the weights introduced for the rain layer and mask layer are denoted as follows: weight $\beta$ for the rain layer, weight $\gamma$ for the mask layer, attenuation coefficient $\lambda$ and atmospheric light intensity $A$ are set to $\beta=1.0$, $\gamma=0.2$, $\lambda=0.025$ and $A=0.5$, respectively. Faced with various atmospheric lighting conditions and different levels of rainfall leading to diverse weather scenarios, modifying the corresponding parameters allows for the synthesis of rainy lane images in the target environment. This capability holds practical significance for depth estimation in complex rainy environments.
During the training of DLKPN, for the first layer KPN network, we set the batch size to 16 and trained for a total of 1000 epochs. We used the inference results of the pre-trained network from the first layer as the training input for the second layer KPN, focusing on removing the rain streak artifacts introduced by the first layer model. With the same batch size, we recorded the network’s performance on the test set in terms of PSNR and SSIM every 100 epochs. After experimentation, we selected the model trained for 400 epochs as the pre-trained network for the second layer of KPN.
III-C Results
TABLE I: Comparison with Current Image Reconstruction Models for Complex Rainy Weather Images
Quantitative Evaluation Metric
Method SSIM PSNR Times (ms)
DRSformer[14] 0.708 22.381 754.72
PreNet[15] 0.685 16.491 95.35
PReNet_r[15] 0.680 16.537 94.35
PRN[15] 0.684 16.440 33.05
PRN_r[15] 0.681 16.396 32.85
DLKPN 0.941 30.601 152.89
TABLE II: Comparison with Depth Estimation Baseline Model in Complex Rainy Conditions
Error Metrics (Lower is better) Accuracy (Higher is better)
Method Abs Rel Sq Rel RMSE RMSE log $\delta<1.25^{1}$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
LapDepth[16] 0.172 1.383 6.304 0.298 0.708 0.846 0.921
KPNDepth(S) 0.110 0.715 4.276 0.202 0.838 0.930 0.981
KPNDepth(D) 0.066 0.447 2.897 0.136 0.941 0.977 0.988
Figure 6: Comparison of Depth Maps Output by Depth Estimation Models on RainKITTI.
III-C1 Image Reconstruction on RainKITTI
The relevant experimental results are shown in Table I. Since the DLKPN method proposed in this paper is based on our construction of the RainKITTI dataset for training and testing, it is not possible to obtain usable prior data from the literature. Moreover, related studies on the fog effects of rain curtains in rain image reconstruction, represented by [17], have not been open-sourced. To evaluate the actual effectiveness of the DLKPN network, we selected the state-of-the-art (SOTA) model DRSformer[14] and the classic lightweight rain removal model PreNet[15] in the field of rain image restoration. After training the above models on the RainKITTI dataset, we compared them with DLKPN in terms of Quantitative Evaluation Metric and inference latency.
The models PreNet, PReNet_r, PRN, and PRN_r in the table are all model architectures proposed in [15], differing in whether recursive layers and intra-level recursive calculations are introduced. The inference latency in Table I is represented by Times(ms), reflecting the inference time of a single image.
From the results, it can be seen that the DLKPN network achieved better SSIM and PSNR on the RainKITTI dataset than the current SOTA method [14] because it considers the darkening of the environment and local fog effects caused by rain when removing rain streaks. By comparing the restored images output by each model in Table 1, it can be observed that the reason for the lower SSIM and PSNR of methods like DRSformer is that these methods still retain some mask and fog information in the reconstructed images. In terms of inference latency, the single-image inference time of DLKPN is 152.89ms. Although there is still a gap compared to the lightweight model PRN_r, it has improved the speed by 4.94 times compared to the SOTA model DRSformer. Considering the SSIM and PSNR performance of the overall model, we believe that the inference latency of DLKPN is acceptable.
III-C2 Depth Estimation of Lane in Rainy Days
The relevant experimental results are shown in Table II, , and the depth estimation results are displayed in Fig. 6. Due to the scarcity of related works on depth estimation in complex environments, there is a lack of existing models and results for comparison. Inspired by similar work [20], we will focus on analyzing the improvement of KPNDepth compared to LapDepth, which serves as the baseline model.
In Table II KPNDepth(S) indicates the depth estimation with only a single-layer pre-trained KPN introduced, while KPNDepth(D) represents using pre-trained DLKPN for prior image reconstruction during depth estimation. According to the results provided in the table, we observe that the LapDepth network with the introduction of pre-trained DLKPN achieves optimal results in both model loss and accuracy. Furthermore, from the depth prediction image in Fig 6 it is evident that the object edges predicted by KPNDepth(D) are clearer and more detailed. During the experiments, we found that due to the influence of rain curtains, the overall depth estimation values of rainy images tend to be erroneously inflated. By comparing the predicted results in Fig. 6, we notice that KPNDepth(D) effectively avoids this issue among the three models.
TABLE III: Results of a Quantitative Study of Ablation on RainKITTI
Method SSIM PSNR Times (ms)
EfDeRain[18] 0.699 21.999 68.54
DLKPN 0.941 30.601 152.89
III-C3 Ablation Study
In the establishment of the DLKPN network, we followed the idea of predicting convolutional kernels instead of directly predicting the pixel values of reconstructed images, as seen in [13][18][19]. Building upon the single KPN network model in [18], we introduced another set of independent pixel-wise convolutional kernels focused on removing rain streak artifacts.
To analyze the impact of the newly introduced kernel set on image reconstruction, we designed a series of ablations to evaluate our method, with quantitative analysis results presented in Table III. The experiments demonstrate that the two sets of independent pixel-wise convolutional kernels introduced by DLKPN significantly enhance the quality of reconstructed images, providing better robustness on lane images in complex weather conditions. However, it is important to note that due to the additional convolutional computation introduced by this network, the inference time for a single image will correspondingly increase.
IV CONCLUSION
In this paper, we address the issue of lane image depth estimation in highly complex rainy environments, which is of great significance for expanding the application scope of depth estimation models and enhancing their robustness in complex lane scenarios.
Building upon an analysis of the limitations of existing rainy weather datasets, we utilized the RCFane algorithm to create the synthetic dataset RainKITTI, capable of reflecting real rainy scenes and rain fog effects. Additionally, a new pre-trained DLKPN network for rainy image reconstruction is proposed in this paper. Integrated with traditional depth estimation models, the KPNDepth depth estimation framework is developed, employing a pixel-wise convolutional approach to restore lost depth information and address rain streak artifacts. This framework achieves more effective image reconstruction in complex rainy lane environments compared to existing methods [14][15], and outperforms the baseline model [16] in terms of accuracy in depth estimation.
Furthermore, the DLKPN framework proposed in this paper is trained on offline data, and experiments indicate its advantage in inference latency, thereby demonstrating good transferability. It offers a feasible approach for depth estimation in complex rainy environments. However, there are still some shortcomings to be addressed. For instance, our current work does not consider nighttime driving scenes, posing a research challenge on how to reconstruct the depth information of low-light objects in such scenarios.